The volume of a gas in a container at a constant temperature varies inversely as the pressure. The volume is 26 cubic centimeters at a pressure of 11 pounds. Use a proportion to find the pressure when the volume is 43 cubic centimeters.
Found 2 solutions by ikleyn, ewatrrr:
Answer by ikleyn(52780)

.

Write a proportion according to the texts' instructions.

Then solve the proportion.

---------------

If you need to learn about proportions,  read the lessons
    - Proportions
    - Using proportions to solve word problems
    - Using proportions to solve word problems in Physics
    - Using proportions to solve Chemistry problems
    - Typical problems on proportions
    - Using proportions to estimate the number of fish in a lake
    - HOW TO algebreze and solve these problems using proportions
    - Problems on proportions for mental solution
    - Selected problems on proportions from the archive
    - OVERVIEW of lessons on proportions
in this site.

These lessons were written specially for you.

Read them and become an expert in this area.

Start from the very first lesson.

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.


--------------

To tutor @ewatrrr

    Please look into the last line in your post source code.


    It's me who added this line to your code.


    May I ask you PLEASE to provide this line (such line) at the end of each your posts ?


    It serves to CLOSE  the font options  of your source code.


    Otherwise, these options extend their influence to the posts of other tutors to the same subject.


    THANK YOU.


    @ikleyn



Answer by ewatrrr(24785)

 
Hi,
The volume of a gas in a container at a constant temperature varies inversely as the pressure.
V = k/P
The volume is 26 cubic centimeters at a pressure of 11 pounds.
26 = k/11
26*11 = k = 286
when the volume is 43 cubic centimeters, find pressure (it will be less as Volume increases)
P = k/V = 286/43 lbs
(Answer is between 6 and 7)
Will Let You finish the exact arithmetic with Your Calculator..
Rounding off Your answer, according to instructions(not given here)
Wish You the Best in your Studies.